McKnight Company is considering two different, mutually exclusive capital expenditure proposals. Profect A will cost $508,824, has an expected useful life of 12 years, a salvage value of zero, and is expected to increase net annual cash flows by $74,100. Project B will cost $341,779, has an expected useful life of 12 years, a salvage value of zero, and is expected to increase net annual cash flows by $50,700. A discount rate of 8% is appropriate for both projects. Compute the net present value and profitability index of each project. (If the net present value is negative, use either a negative sign preceding the number eg-45 or parentheses eg (45). Round present value answers to 0 decimal places, e.g. 125 and profitability index answers to 2 decimal places, e.g. 15.25. For calculation purposes, use 5 decimal places as displayed in the factor table provided.) Net present value - Project A s Profitability index Project A et present value-Project B Profitability index Project B Which project should be accepted based on Net Present Value? Project A should be accepted. Which project should be accepted based on profitability index? Project B should be accepted.
Answer #1
MC Knight Company
P.v of annuity factor 8% for 12 years=(A) 7.5361
Project A Project B
Cost=(B) 508824 341779
Expected increase in net annual cash flow=(C ) 74100 50700
P.V of cash flow=(D )=(A)*(C) $       5,58,425 $       3,82,080
NPV=Present value of inflow-Present value of outflow
NPV=(D)-(B) 49601 40301
Project A should be accepted on the basis of NPV
Profitability Index=Present value of cash flow/Present value of cash outflow
P.V of cash flow=(A) 558425.01 382080.27
P.V of cash outflow=(B) 508824 341779
Profitability Index=(A)/(B) 1.10 1.12